# Rationality based Innate-Values-driven Reinforcement Learning

### v1

@@ 10             Set
$y_{j}=\left\{\begin{array}[]{lr}r_{j},~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}\text{for terminal }\phi_{j+1};\\ r_{j}+\gamma\max_{w,a}Q(\phi_{j+1},w^{\prime},a^{\prime};\theta),~{}~{}~{}\text{else}.\end{array}\right.$
11             Perform a gradient descent step on $(y_{j}-Q(\phi_{j},w_{j},a_{j};\theta))^{2}$ according to Eq. (14).
@@
Algorithm 1 Innate-Values-driven RL (IVRL) DQN Version
Moreover, we utilize the experience replay technique lin1992reinforcement , which stores the agent’s experiences at each time-step, $e_{t}=(s_{t},w_{t},a_{t},r_{t},s_{t+1})$ in a data-set $\mathcal{D}=e_{1},\dots,e_{N}$, pooled over many episodes into a replay memory. During the algorithm’s inner loop, we apply Q-learning updates, or minibatch updates, to samples of experience, $e\sim\mathcal{D}$, drawn at random from the pool of stored samples. After performing experience replay, the agent selects and executes an action according to an $\epsilon$-greedy policy, as we discussed. Since implementing arbitrary length histories as inputs to a neural network is difficult, we use a function $\phi$ to produce our action-innate-value Q-function. Alg. 1 presents the algorithm of the IVRL DQN.
2.6 IVRL Advantage Actor-Critic (A2C) Model
Figure 6: The architecture of the IVRL Actor-Critic Version.
Furthermore, we extend our IVRL method to the Advantage Actor-Critic (A2C) version. Specifically, our IVRL A2C maintains a policy network $\pi(a_{t}|s_{t};\theta)$, a needs network $\omega(w_{t}|s_{t};\delta)$, and a utility value network $u(s_{t},a_{t};\varphi)$. Since the reward in each step is equal to the current utilities $u(s_{t},a_{t})$ multiplying the corresponding weight of needs, the state innated-values function can be approximated by presenting it as Eq. (15). Then, we can get the policy gradient Eq. (16) and needs gradient Eq. (17) of the Eq. (15) deriving $V(s;\theta,\delta)$ according to the Multi-variable Chain Rule, respectively. We can update the policy network $\theta$ and needs network $\delta$ by implementing policy gradient and needs gradient, and using the temporal difference (TD) to update the value network $\varphi$.
@@ 14             Optimize $U_{\varphi}$ using $\nabla L(\varphi)$
Algorithm 2 Innate-Values-driven RL (IVRL) Advantage Actor-Critic (A2C) Version
$\begin{split}V_{\pi,\omega}(s)&=\sum_{a_{t},w_{t}}\pi(a_{t}|s_{t})\cdot\omega(w_{t}|s_{t})\cdot u(s_{t},a_{t})\\ &\approx\sum_{a,w}\pi(a_{t}|s_{t};\theta)\cdot\omega(w_{t}|s_{t};\delta)\cdot u(s_{t},a_{t};\varphi)=V(s;\theta,\delta,\varphi)\end{split}$ (15)
grad $\begin{split}\text{grad~{}}V(a_{t},\theta_{t})=V_{\theta}(s;\theta,\delta,\varphi)=\frac{\partial V(s;\theta,\delta,\varphi)}{\partial\theta}=\frac{\partial\pi(a_{t}|s_{t};\theta)}{\partial\theta}\cdot\omega(w_{t}|s_{t};\delta)\cdot u(s_{t},a_{t};\varphi)\end{split}$ (16)
grad $\begin{split}\text{grad~{}}V(w_{t},\delta_{t})=V_{\delta}(s;\theta,\delta,\varphi)=\frac{\partial V(s;\theta,\delta,\varphi)}{\partial\delta}=\pi(a_{t}|s_{t};\theta)\cdot\frac{\partial\omega(w_{t}|s_{t};\delta)}{\partial\delta}\cdot u(s_{t},a_{t};\varphi)\end{split}$ (17)
Using an estimate of the utility $u$ function as the baseline function, we subtract the V value term as the advantage value. Intuitively, this means how much better it is to take a specific action and a needs weight compared to the average, general action and the needs weights at the given state Eq. (18). Fig. 6 illustrates the architecture of the IVRL actor-critic version and Alg. 2 presents the algorithm of the IVRL A2C.
$A(s_{t},a_{t})=U(s_{t},a_{t})-V(s_{t})$ (18)
3 Experiments
@@
Considering cross-platform support and the ease of creating custom scenarios, we selected the VIZDoom Role-Playing Game (RPG) reinforcement learning test platform Kempka2016ViZDoom ; Wydmuch2019ViZdoom to evaluate the performance of the proposed innate-value-driven reinforcement learning model. We choose four scenarios: Defend the Center, Defend the Line, Deadly Corridor, and Arens (Fig. 7), and compare our models with several benchmark algorithms, such as DQN mnih2015human , DDQN wang2016dueling , A2C mnih2016asynchronous , and PPO schulman2017proximal . These models were trained on an NVIDIA GeForce RTX 3080Ti GPU with 16 GiB of RAM.
3.1 Environment Setting
In our experiments, we define four categories of utilities (health points, amount of ammo, environment rewards, and number of killed enemies), presenting three different levels of needs: low-level safety and basic needs, medium-level recognition needs, and high-level achievement needs. When the agent executes an action, it receives all the corresponding innate utilities, such as health points and ammo costs, and external utilities, such as environment rewards (living time) and the number of killed enemies. At each step, the agent can calculate the rewards for the action by multiplying the current utilities and the needed weight for them. In our experiments, the initial needs weight for each utility category is 0.25, which has been fixed in the benchmark DRL algorithms’ training, such as DQN, DDQN, and PPO. For more details about the experiment code, please check the supplementary materials.
@@ 3.1 Environment Setting
3.2 Evaluation
Figure 8: The performance comparison of IVRL DQN and A2C agents with DQN, DDQN, PPO, and A2C in the VIZDoom.
@@ 3.2 Evaluation
The performance of the proposed IVRL DQN and A2C models is shown in the Fig 8. Fig. 7(a), 7(b), 7(c), and 7(d) demonstrate that IVRL models can achieve higher average scores than traditional RL benchmark methods (Fig. 8, 8, 8, and 8). Especially for the IVRL A2C algorithm, it presents more robust, stable, and efficient properties than the IVRL DQN model. Although the IVRL DQN shows better performance in the Arena scenario (Fig. 8), the small perturbation introduced by the innate-values utilities may have made the network weights in some topology difficult to reach convergence.
@@
Furthermore, we also analyze their corresponding tendencies in different scenarios to compare the needs weight differences between the IVRL DQN and A2C models. In the defend-the-center and defend-the-line experiments, each category of the need weight in the IVRL DQN model does not split and converges to a specific range compared with its initial setting in our training (Fig. 8 and 8). In contrast, the weights of health depletion, ammo cost, and sub-goal (environment rewards) shrink to approximately zero, and the weight of the number of killed enemies converges to one in the IVRL A2C model. This means that the top priority of the IVRL A2C agent is to eliminate all the threats or adversaries in those scenarios so that it can survive, which is similar to the Arena task. According to the performance in those three scenarios (Fig. 8, 8, and 8), the IVRL A2C agent represents the characteristics of bravery and fearlessness, much like the human hero in a real battle. However, in the deadly corridor mission, the needs weight of the task goal (getting the vest) becomes the main priority, and the killing enemy weight switches to the second for the IVRL A2C agent (Fig. 8). They converge to around 0.6 and 0.4, respectively. In training, by adjusting its different needs weights to maximize rewards, the IVRL A2C agent develops various strategies and skills to kill the encounter adversaries and get the vast efficiently, much like a military spy.
@@
In our experiments, we found that selecting the suitable utilities to consist of the agent innate-values system is critically important for building its reward mechanism, which decides the training speed and sample efficiency. Moreover, the difference in the selected utility might cause some irrelevant experiences to disrupt the learning process, and this perturbation leads to high oscillations of both innate-value rewards and needs weight.
@@
Generally, the innate value system serves as a unique reward mechanism driving agents to develop diverse actions or strategies satisfying their various needs in the systems. It also builds different personalities and characteristics of agents in their interaction. From the environmental perspective, due to the various properties of the tasks, agents need to adjust their innate value system (needs weights) to adapt to different tasks’ requirements. These experiences also shape their intrinsic values in the long term, similar to humans building value systems in their lives. Moreover, organizing agents with similar interests and innate values in the mission can optimize the group utilities and reduce costs effectively, just like “Birds of a feather flock together." in human society.
4 Conclusion
@@
This paper introduces the innate-values-driven reinforcement learning (IVRL) model mimicking the complex behaviors of agent interactions. By adjusting needs weights in its innate-values system, it can adapt to different tasks representing corresponding characteristics to maximize the rewards efficiently. For theoretical derivation, we formulated the IVRL model and proposed two types of IVRL models: DQN and A2C. Furthermore, we compared them with benchmark algorithms such as DQN, DDQN, A2C, and PPO in the RPG reinforcement learning test platform VIZDoom. The results prove that rationally organizing various individual needs can effectively achieve better performance.
@@
For future work, we want to improve the IVRL further and develop a more comprehensive system to personalize individual characteristics to achieve various tasks testing in several standard MAS testbeds, such as StarCraft II, OpenAI Gym, Unity, etc. Especially in multi-object and multi-agent interaction scenarios, building the awareness of AI agents to balance the group utilities and system costs and satisfy group members’ needs in their cooperation is a crucial problem for individuals learning to support their community and integrate human society in the long term. Moreover, integrating efficient deep RL algorithms with the IVRL can help agents evolve diverse skills to adapt to complex environments in MAS cooperation. Furthermore, implementing the IVRL in real-world systems, such as human-robot interaction, multi-robot systems, and self-driving cars, would be challenging and exciting.

### v2

@@ 10             Set
$y_{j}=\left\{\begin{array}[]{lr}r_{j},~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}\text{for terminal }\phi_{j+1};\\ r_{j}+\gamma\max_{w,a}Q(\phi_{j+1},w^{\prime},a^{\prime};\theta),~{}~{}~{}\text{else}.\end{array}\right.$
11             Perform a gradient descent step on $(y_{j}-Q(\phi_{j},w_{j},a_{j};\theta))^{2}$ according to (16).
@@
Algorithm 1 Innate-Values-driven DQN (IV-DQN)
Moreover, we utilize the experience replay technique [31], which stores the agent’s experiences at each time-step, $e_{t}=(s_{t},w_{t},a_{t},r_{t},s_{t+1})$ in a data-set $\mathcal{D}=e_{1},\dots,e_{N}$, pooled over many episodes into a replay memory. During the algorithm’s inner loop, we apply Q-learning updates, or minibatch updates, to samples of experience, $e\sim\mathcal{D}$, drawn at random from the pool of stored samples. After performing experience replay, the agent selects and executes an action according to an $\epsilon$-greedy policy, as we discussed. Since implementing arbitrary length histories as inputs to a neural network is difficult, we use a function $\phi$ to produce our action-innate-value Q-function. Alg. 1 presents the algorithm of the IV-DQN.
II-F IVRL Advantage Actor-Critic (A2C) Model
Furthermore, we extend our IVRL method to the Advantage Actor-Critic (A2C) version. Specifically, our IV-A2C maintains a policy network $\pi(a_{t}|s_{t};\theta)$, a needs network $\omega(w_{t}|s_{t};\delta)$, and a utility value network $u(s_{t},a_{t};\varphi)$. Since the reward in each step is equal to the current utilities $u(s_{t},a_{t})$ multiplying the corresponding weight of needs, the state innated-values function can be approximated by presenting it as (17). Then, we can get the policy gradient (18) and needs gradient (19) of the (17) deriving $V(s;\theta,\delta)$ according to the Multi-variable Chain Rule, respectively. We can update the policy network $\theta$ and needs network $\delta$ by implementing policy gradient and needs gradient, and using the temporal difference (TD) to update the value network $\varphi$.
Figure 6: The architecture of the IVRL Actor-Critic model.
@@ 14             Optimize $U_{\varphi}$ using $\nabla L(\varphi)$
Algorithm 2 Innate-Values-driven Advantage Actor-Critic (IV-A2C)
@@
$\begin{split}V_{\pi,\omega}(s)&=\sum_{a_{t},w_{t}}\pi(a_{t}|s_{t})\cdot\omega(w_{t}|s_{t})\cdot u(s_{t},a_{t})\\ &\approx\sum_{a,w}\pi(a_{t}|s_{t};\theta)\cdot\omega(w_{t}|s_{t};\delta)\cdot u(s_{t},a_{t};\varphi)\\ &=V(s;\theta,\delta,\varphi)\end{split}$ (17)
Using an estimate of the utility $u$ function as the baseline function, we subtract the V value term as the advantage value. Intuitively, this means how much better it is to take a specific action and a needs weight compared to the average, general action and the needs weights at the given state (20). Fig. 6 illustrates the architecture of the IVRL actor-critic version and Alg. 2 presents the algorithm of the IV-A2C.
grad $\begin{split}\text{grad~{}}V(a_{t},\theta_{t})&=V_{\theta}(s;\theta,\delta,\varphi)=\frac{\partial V(s;\theta,\delta,\varphi)}{\partial\theta}\\ &=\frac{\partial\pi(a_{t}|s_{t};\theta)}{\partial\theta}\cdot\omega(w_{t}|s_{t};\delta)\cdot u(s_{t},a_{t};\varphi)\end{split}$ (18)
grad $\begin{split}\text{grad~{}}V(w_{t},\delta_{t})&=V_{\delta}(s;\theta,\delta,\varphi)=\frac{\partial V(s;\theta,\delta,\varphi)}{\partial\delta}\\ &=\pi(a_{t}|s_{t};\theta)\cdot\frac{\partial\omega(w_{t}|s_{t};\delta)}{\partial\delta}\cdot u(s_{t},a_{t};\varphi)\end{split}$ (19)
$A(s_{t},a_{t})=U(s_{t},a_{t})-V(s_{t})$ (20)
III Experiments
Figure 8: The performance comparison of IV-DQN and IV-A2C agents with DQN, DDQN, PPO, and A2C in the VIZDoom.
Since the IVRL model differs from the traditional RL model, it uses need weights and corresponding utilities from the internal and external environment to calculate innate value rewards. The traditional environment feedback-based reward RL platform can not be used in the IVRL experiments. Considering that the VIZDoom testbed [26, 27] can customize the experiment environment and define various utilities based on different tasks and cross-platform, we selected it to evaluate evaluate our IVRL model. We choose four scenarios: Defend the Center, Defend the Line, Deadly Corridor, and Arens (Fig. 7), and compare our models with several benchmark algorithms, such as DQN [22], DDQN [23], A2C [24], and PPO [25]. These models were trained on an NVIDIA GeForce RTX 3080Ti GPU with 16 GiB of RAM. Moreover, we open the source of the IVRL on GitHub: https://github.com/is3rlab/Innate-Values-driven-Reinforcement-Learning.
III-A Environment Setting
In our experiments, we define four categories of utilities (health points, amount of ammo, environment rewards, and number of killed enemies), presenting three different levels of needs: low-level safety and basic needs, medium-level recognition needs, and high-level achievement needs. When the agent executes an action, it receives all the corresponding innate utilities, such as health points and ammo costs, and external utilities, such as environment rewards (living time) and the number of killed enemies. At each step, the agent can calculate the rewards for the action by multiplying the current utilities and the needed weight for them. In our experiments, we assume that the agent has no bias or preference at the beginning of the game. Therefore, the initial needs weight for each utility category is 0.25, fixed in the benchmark DRL algorithms’ training, such as DQN, DDQN, and PPO. For more details about the experiment code, please check the supplementary materials.
@@ III-A Environment Setting
III-B Evaluation
@@
The performance of the proposed IV-DQN and IV-A2C models is shown in the Fig. 8, 8, 8, and 8 demonstrate that IVRL models can achieve higher average scores than traditional RL benchmark methods. Especially for the IV-A2C algorithm, it presents more robust, stable, and efficient properties than other models.
@@
Moreover, we analyze their corresponding tendencies in different scenarios to compare the needs weight differences between the IV-DQN and IV-A2C models. In the defend-the-center and defend-the-line experiments, each category of the need weight in the IV-DQN model does not split and converges to a specific range compared with its initial setting in our training (Fig. 8 and 8). In contrast, the weights of health depletion, ammo cost, and sub-goal (environment rewards) shrink to approximately zero, and the weight of the number of killed enemies converges to one in the IV-A2C model. This means that the top priority of the IV-A2C agent is to eliminate all the threats or adversaries in those scenarios so that it can survive, which is similar to the Arena task. According to the performance in those three scenarios (Fig. 8, 8, and 8), the IV-A2C agent represents the characteristics of bravery and fearlessness, much like the human hero in a real battle. However, in the deadly corridor mission, the needs weight of the task goal (getting the vest) becomes the main priority, and the killing enemy weight switches to the second for the IV-A2C agent (Fig. 8). They converge to around 0.6 and 0.4, respectively. In training, by adjusting its different needs weights to maximize rewards, the IV-A2C agent develops various strategies and skills to kill the encounter adversaries and get the vast efficiently, much like a military spy.
@@
In our experiments, we found that selecting the suitable utilities to consist of the agent innate-values system is critically important for building its reward mechanism, which decides the training speed and sample efficiency. Moreover, the difference in the selected utility might cause some irrelevant experiences to disrupt the learning process, and this perturbation leads to high oscillations of both innate-value rewards and needs weight. Furthermore, the IV-DQN performs better in the Arena than any other algorithm (Fig. 8). However, in other experiments, the IV-A2C’s performance is better than the IV-DQN. It reflects that, due to different task scenarios, the small perturbation introduced by the innate-values utilities may have made it difficult for the network weights in some topologies to reach convergence. Generally speaking, the performances of IV-DQN and IV-A2C are generally better than traditional A2C, DQN, and PPO.
@@
The innate value system serves as a unique reward mechanism driving agents to develop diverse actions or strategies satisfying their various needs in the systems. It also builds different personalities and characteristics of agents in their interaction. From the environmental perspective, due to the various properties of the tasks, agents need to adjust their innate value system (needs weights) to adapt to different tasks’ requirements. These experiences also shape their intrinsic values in the long term, similar to humans building value systems in their lives.
IV Conclusion
@@
This paper introduces a new RL model from individual intrinsic motivations perspectives termed innate-values-driven reinforcement learning (IVRL). It is based on the expected utility theory to model mimicking the complex behaviors of agent interactions in its evolution. By adjusting needs weights in its innate-values system, it can adapt to different tasks representing corresponding characteristics to maximize the rewards efficiently. In other words, through interacting with other agents and environments, the agent builds its unique value system to adapt to them, the same as ”Individual is the product of their own environment.” For theoretical derivation, we formulated the IVRL model and proposed two types of IVRL models: IV-DQN and IV-A2C. Furthermore, we compared them with benchmark algorithms in the RPG reinforcement learning test platform VIZDoom. The results prove that rationally organizing various individual needs can effectively achieve better performance. Moreover, in the multi-agent setting, organizing agents with similar interests and innate values in the mission can optimize the group utilities and reduce costs effectively, just like “Birds of a feather flock together.” in human society. Especially combined with AI agents’ capacity to aid decision-making, it will open up new horizons in human-multi-agent collaboration. This potential is crucially essential in the context of interactions between human agents and intelligent agents when considering establishing stable and reliable relationships in their cooperation, particularly in adversarial and rescue mission environments.
@@
For future work, we want to improve the IVRL further and develop a more comprehensive system to personalize individual characteristics to achieve various tasks testing in several standard MAS testbeds, such as StarCraft II, OpenAI Gym, Unity, etc. Especially in multi-object and multi-agent interaction scenarios, building the awareness of AI agents to balance the group utilities and system costs and satisfy group members’ needs in their cooperation is a crucial problem for individuals learning to support their community and integrate human society in the long term. Furthermore, implementing the IVRL in real-world systems, such as human-robot interaction, multi-robot systems, and self-driving cars, would be challenging and exciting.
Acknowledgments
This work is supported by the NSF Foundational Research in Robotics (FRR) Award 2348013.